The energy from radiation can be used to cause the rupture of chemical bonds. A minimum energy of 242 kJ/mol is required to break the chlorine–chlorine bond in Cl2.

Part A

What is the longest wavelength of radiation that possesses the necessary energy to break the bond?

Express your answer with the appropriate units.

λ =

SubmitMy AnswersGive Up

Part B
Answer #1

E = 242 kJ/mol

E = 242000 J/mol --> 1 mol = 6.02"*10^23

E = 4.019*10^-19 J/molecule

relate with:

WL = h c / E

WL = wavelenght

c = speed of light 3*10^8 m/s

h = Planck Constant =   6.626 x 10 - 34 J s

Susbtitute all data:

WL = h c / E = ( 6.626 x 10 - 34 J s ) (3*10^8 m/s)/ (4.019*10^-19 J/molecule)

WL = 4.94*10^-7 m

WL = 494 nm